OpenStudy (anonymous):

3^16/3^3 Divide and simplify.

OpenStudy (anonymous):

3^16-3 3^13

OpenStudy (anonymous):

That's what I thought, but I wasn't sure. Thank you. :)

OpenStudy (anonymous):

Wouldn't I divide 3 by 3 and make it 1^13?

OpenStudy (anonymous):

That part's still got me. Why would I ignore the whole number?

OpenStudy (anonymous):

\[{3^{16} \over 3^{3} } = 3^{16} \times 3^{-3} = 3^{16-3} = 3^{13}\]

OpenStudy (anonymous):

Thank you. :)

OpenStudy (amistre64):

its not a whole number; its a base number :)

OpenStudy (amistre64):

3.3.3.3.3.3.3.3.3.3.3.3.3.3.3.3 --------------------------- 3.3.3 3.3.3.3.3.3.3.3.3.3.3.3.3 = 3^13

OpenStudy (amistre64):

a^4 a.a.a.a --- = ------- = a^1 = a a^3 a.a.a